OpenStudy (anonymous):

[4.06] What is the equation of the line that passes through the point (4, 1) and has a slope of 5? @ghazi

OpenStudy (anonymous):

PLEASECHELP @ghazi

OpenStudy (ghazi):

lets say equation of line is y=mx+c now, y=5x+c is the line put x=4 and y=1 1=5*4+c c=1-20=-19 now y=5x-19 is the line

OpenStudy (ghazi):

@julianmiddleton clear?

OpenStudy (anonymous):

ohh i get it can you help with a couple more @ghazi

OpenStudy (anonymous):

Monisha decides to join a gym for a minimum of one month to train for a triathlon. The gym charges a joining fee of $100 and a monthly fee of $37. If x represents the number of months that Monisha is a member of the gym, the equation below can be used to determine C, her total membership fee for that duration of time. 100 + 37x = C Monisha has allocated a maximum of $285 to spend on her gym membership. Which number line shows the possible number of months that Monisha can be a member of the gym?

OpenStudy (ghazi):

i will definitely help you but OS is running too slow right now, so a bit later :)